What is the potential at a point 18 m away from a 17.14 mC charge?


1
Expert's answer
2022-02-06T14:34:41-0500

Given:

"r=18\:\rm m"

"q=17.14*10^{-3}\:\rm C"


The ptential of electric field

"V=k\frac{q}{r}=9*10^9*\frac{17.14*10^{-3}}{18}=8.57*10^6\:\rm V"